# Efficient and Generalizable Certified Unlearning: A Hessian-free Recollection Approach

### v3

@@ 3 Main Intuition and Solution Justification
where $\mathbf{I}$ denotes identity matrix of appropriate size. Combining the observations in Cases 1 and 2, we now define the recollection matrix $\mathbf{M}_{e,b(u)}$ and affine stochastic recursion approximator $\mathbf{a}_{E,B}^{-u}$, backtracking the overall difference between the retrained model and the learned model:
$\mathbf{a}_{E,B}^{-u}:=\sum_{e=0}^{E}\mathbf{M}_{e,b(u)}\nabla\ell(\mathbf{w}_{e,b(u)};u),\ \mathbf{M}_{e,b(u)}:=\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\prod^{E}_{k=e}\prod_{b=b(u)+1}^{B-1}\Big{(}\mathbf{I}\!-\!\frac{\eta_{k,b}}{|\mathcal{B}_{k,b}|}\mathbf{H}_{k,b}\Big{)}.$ (10)
We will see that $\mathbf{a}_{E,B}^{-u}$ can approximate the impact of sample $u$ on the training trajectory [^1] Similar analysis has been conducted in other lines of work to investigate the behavior of SGD dynamics, e.g. reference [29] focused on the properties of the matrix $\mathbf{I}-\frac{\eta}{|\mathcal{B}|}\mathbf{H}$, which is termed as multiplicative noise, serving as the main source of heavy-tails in SGD dynamics and governing the update behavior of $\mathbf{w}_{e,b}$. as follows,
$\mathbf{w}_{E,B}^{-u}$ $-\mathbf{w}_{E,B}\approx\mathbf{a}_{E,B}^{-u},$ (11)
That is, a simple vector $\mathbf{a}_{E,B}^{-u}$ addition to model $\mathbf{w}_{E,B}$ can obtain an approximate retrained model for forgetting the sample $u$. We define $\Delta_{E,B}^{-u}=\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}-\mathbf{a}_{E,B}^{-u}$ as the approximation error.
Our designed approximator in (10) enables us to exploit the Hessian vector product (HVP) technique [7] to compute it without explicitly computing the Hessian. When computing the product of the Hessian matrix and an arbitrary vector for the deep neural networks, the HVP technique first obtains gradients during the first backpropagation, multiplying the gradients with the vector, and then performing backpropagation again to obtain the results. This results in a time complexity of $\mathcal{O}(d)$. We note that the existing second-order methods require to proactively store a Hessian matrix (or its inverse) [2, 3, 4]. It is thus unable to deploy the HVP in their settings.
Note that our proposed analysis does not require to invert Hessian matrix, distinguishing it from existing Newton-like methods [2, 3, 4]. This eliminates the need for assuming strong convexity in order to ensure the invertibility of the Hessian. In our analysis, we consider the approximation error under general non-convex non-smooth settings (e.g., deep neural networks) by examining the spectral radius $\rho$ of the recollection matrix $\mathbf{M}_{e,b(u)}$, which characterizes the discrepancy between the retrained and learned model $\mathbf{w}_{E,B}^{-U}\!-\!\mathbf{w}_{E,B}$. Moreover, when it comes to forgetting a set of samples, we have,
Theorem 1 (Additivity).
When $m$ sequential deletion requests arrive, the sum of $m$ approximators is equivalent to performing batch deletion simultaneously, i.e., for $u_{1},\ldots,u_{m}$ sequence of continuously arriving deletion requests, we demonstrate that $\mathbf{a}_{E,B}^{-\sum_{j=1}^{m}u_{j}}=\sum_{j=1}^{m}\mathbf{a}_{E,B}^{-u_{j}}$.
This elucidates that by storing the impact of updates for each sample, nearly instantaneous data removal can be achieved through a series of vector additions.
4 Theoretical Analysis and Algorithm Design
In this section, we first analyze the approximation error $\Delta_{E,B}^{-U}$ under non-convex non-smooth setting. We then demonstrate that under the same assumptions, our method can almost obtain an identical retrained model and achieve better unlearning guarantee and generalization results compared to previous theoretical unlearning works [2, 3, 4]. We further develop our proposed unlearning algorithm, which is one of the most efficient and nearly instantaneous. Finally, we provide storage/computation complexity analysis of proposed algorithm and show that it is more efficient compared to prior works.
4.1 Approximation Error Analysis
To analyze the approximation error $\Delta_{E,B}^{-U}$, we introduce the following lemmata. We first bound the approximation error through gradient clipping, which is widely adopted to achieve favorable performances in deep network training, by ensuring gradient norms are not excessively large [30]. We further show that the gradient clipping with geometrically decaying stepsizes also facilitates us to derive an upper bound for the remainder term ${o(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})}$ in the following:
Lemma 2.
In the learning stage, we threshold the stochastic gradient norm at $C$, i.e., $\|\nabla\ell(\mathbf{w};z)\|\!\!\leq\!C$. To simplify the expression, we define that SGD performs a total of $t\!\in\!\!\{0,\ldots,T\}$ steps for $\mathbf{w}_{e,b}$, where $t\!=\!eB+b$. Consider geometrically decaying stepsizes satisfying $\eta_{t+1}\!\!=\!q\eta_{t}$, where $10\!<q\!<$ 1 and $\eta=\eta_{0}$. Therefore, after $t$ steps, for any forgetting set $U$, we have:
$\|\mathbf{w}_{e,b}^{-U}-\mathbf{w}_{e,b}\|\leq 2\eta C\frac{1-q^{t}}{1-q}.$ (12)
See proof in Subsection B.3. Then we further analyze the properties of the Hessian matrix:
Lemma 3.
For all $e,b$ and vector $\mathbf{v}\in\mathbb{R}^{d}$, assume that exists $\lambda$ and $M$ such that $\lambda\mathbf{I}\preceq\nabla^{2}\ell(\mathbf{w}_{e,b};z_{i})\preceq M\mathbf{I}$, and let $\rho=\text{max}\{|1-\eta_{e,b}\lambda|,|1-\eta_{e,b}M|\}$ be the spectral radius of $\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$, we have $\|(\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b})\mathbf{v}\|\leq\rho\|\mathbf{v}\|.$
Therefore, combining (10), Lemma 2 and Lemma 3, we have the following upper bound:
Theorem 4 (Unlearning Guarantee).
For any subset of samples $U$=$\{u_{j}\}_{j=1}^{m}$ to be forgotten, the unlearned model based on the approximator $\mathbf{a}_{E,B}^{-U}$ in (10) leads to an approximate error bounded by
$\|\Delta_{E,B}^{-U}\|=\|\mathbf{w}_{E,B}^{-U}-\mathbf{w}_{E,B}-\mathbf{a}_{E,B}^{-U}\|\leq 2\eta C\zeta_{T}^{-U},$ (13)
where $T$ is the total number of update steps, and we define $\zeta_{T}^{-U}=\mathcal{O}(\frac{\rho^{T}-q^{T}}{\rho-q}-\frac{\rho^{T}-q^{2T}}{\rho-q^{2}}+\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}})$.
See proof in Subsection B.4. Intuitively, Theorem 4 shows that through stochastic matrix recursions of (8) and (9), the remainder terms $o(\mathbf{w}_{e,b}^{-u}\!\!-\!\!\mathbf{w}_{e,b})$ of the previous rounds are scaled by subsequent recollection matrices $\mathbf{M}_{e,b}$, and generate a new remainder term in this round, which is a polynomial multiply geometric progression. As a result, when $\rho\!<\!1$ (e.g., Assumption 1), in a scenario where $T$ is sufficiently large, the error tends towards zero. This implies that the unlearned model can nearly approximate the retrained model. Notably, it is worth mentioning that the upper bounds provided by previous theoretical works [2, 3, 4] are non-zero under Assumption 1.
4.2 Generalization Properties Analysis
To discuss the performance guarantee of our proposed method, we further analyze the generalization performance. We present a generalization of the excess risk as defined in problem (1). Subsequently, we conduct an asymptotic analysis of our method, showing that for strongly convex and smooth objectives, our approach can asymptotically approach the retrained objective.
We adopt the following standard assumptions (e.g., similar to prior certified unlearning works [3, 4]):
Assumption 1.
For any $z$ and $\mathbf{w}$, the loss $\ell(\mathbf{w};z)$ is $\lambda$-strongly convex, $L$-lipschitz and $M$-smooth.
Theorem 5 (Generalization Guarantee).
Let Assumption 1 hold. Choose $\eta\leq\frac{2}{M+\lambda}$, then we have $\rho<1$. Let ${\mathbf{w}}^{*}$ be the minimizer of the population risk in problem (1) and $T$ be the total number of update steps. Considering the forgetting dataset $U$ and an unlearned model $\tilde{\mathbf{w}}_{E,B}^{-U}$, we have following excess risk,

### v4

@@ 4.4 Comparison to Existing Approaches
Infinitesimal Jackknife (IJ). Suriyakumar et al. (2022) builded upon the work of NS and facilitated forgetting by pre-computing and storing the inverse Hessian. The update for NS thus is as follows:
[…]
These methods fail to use HVP, for Hessian $\sum_{j\in U}\nabla^{2}\ell(\hat{\mathbf{w}};u_{j})$ in NS and gradient $\nabla\ell(\hat{\mathbf{w}},u_{j})$ in both NS and IJ, (i) forgetting sample $u_{j}$ is unknown before deletion request arrives, and (ii) the learned empirical risk minimzer $\hat{\mathbf{w}}$ will be deleted after processing a deletion request, while subsequent unlearned model is also unknown as forgetting sample $u_{j}$ is unpredictable. Therefore, explicit pre-computing is required for $\sum_{i=1}^{n}\nabla^{2}\ell(\hat{\mathbf{w}};z_{i})$ in NS and ${(\sum_{i=1}^{n}\nabla^{2}\ell(\hat{\mathbf{w}};z_{i}))}^{-1}$ in IJ.
Comparison of Precomputation. For a single data, Algorithm 1 entails computing the HVP $\mathcal{O}{({E}n/{|\mathcal{B}|})}$ times. To precompute the Hessian-free (HF) approximators $\mathbf{a}^{-u}_{\text{HF}}$ for all $n$ samples, the complexity is $\mathcal{O}(n^{2}d{E}/{|\mathcal{B}|})$. Typically, the ratio of $E$ to $|\mathcal{B}|$ is much smaller than $d$ or $n$ (Smith (2018)), which simplifies the complexity to $\mathcal{O}(n^{2}d)$. HF method is highly efficient when dealing with overparameterized models, especially when $d\!\gg\!n$ [^4] Deep networks for computation vision tasks are typically trained with a number of model parameters that far exceeds the size of the training dataset (Chang et al. (2021)). For general neural language model, the empirical data-to-parameter scaling law is also expected to remain at $n\sim d^{0.74}$, as suggested in Kaplan et al. (2020). This is because computing the full Hessian for NS requires $\mathcal{O}(nd^{2})$ time, and computing and inverting the Hessian $\mathbf{H}^{-1}_{\text{IJ}}$ for IJ requires $\mathcal{O}(d^{3}\!+\!nd^{2})$.
Comparison of Storage. When we compute the approximator, it requires storing $\mathcal{O}{({E}n/{|\mathcal{B}|})}$ models in Stage I, leading to a maximum complexity of $\mathcal{O}(nd{E}/{|\mathcal{B}|}+nd)$, which can simplify to $\mathcal{O}(2nd)$. After complete precomputing, Algorithm 1 only needs the storage of $\mathcal{O}(nd)$. This is lower than the overhead introduced by storing matrices for NS and IJ in overparameterized models, which is $\mathcal{O}(d^{2})$.
Comparison of Unlearning Computation. When an unlearning request with $m$ forgetting data arrives, Algorithm 1 utilizes precomputed and prestored approximators to achieve forgetting, requiring a simple vector addition that takes $\mathcal{O}(md)$ time. For NS, it requires computing and inverting different Hessian that depends on the user requesting the deletion. This necessitates substantial computational cost of $\mathcal{O}(d^{3}\!+\!md^{2}\!+\!md)$ each time a deletion request arrives. Although IJ reduces this to $\mathcal{O}(d^{2}\!+\!md)$ by precomputing the inverse Hessian, it is still not as efficient compared to our method.
5 Experiments from Developer’s Perspective
We devised the following experimental components: ❶ Verification and ❷ Application. Verification experiments are centered on evaluating the disparity between the unlearning model $\bar{\mathbf{w}}_{E,B}^{-U}$ and retrained models ${\mathbf{w}}_{E,B}^{-U}$, across both convex and non-convex settings. Application experiments are geared towards evaluating the performance of different unlearning algorithms after perturbing $\tilde{\mathbf{w}}_{E,B}^{-U}$ with noise $\mathbf{n}$ with a specific focus on determining the accuracy, pre-computation/storage, and unlearning time. Considering the space limit, we defer a portion of experiments to the Appendix D.
5.1 Verification Experiments
Our principal aim is to validate the differences between the proposed Hessian-free (HF) approximator $\mathbf{a}^{-U}_{\text{HF}}$ and the actual $\mathbf{w}_{E,B}^{-U}\!-\!\mathbf{w}_{E,B}$, where a smaller difference indicates the unlearned model is closer to retrained model and also reflects the degree of data forgetting. Therefore, we evaluate the degree of forgetting from two perspectives: distance and correlation. Specifically, we use the $L^{2}$-norm $\|\mathbf{w}_{E,B}^{-U}\!\!-\!\mathbf{w}_{E,B}\!\!-\!\mathbf{a}^{\!-U}\|$ to measure the distance metric. A smaller distance metric indicates the approximators accurately capture the disparity between the retrained and learned models. Additionally, we utilize the Pearson (Wright (1921)) and Spearman (Spearman (1961)) coefficients to assess the correlation between $\mathbf{a}^{-U}$ and $\mathbf{w}_{\!E,B}^{-U}\!-\!\mathbf{w}_{\!E,B}$ by mapping them from high-dimensional space to scalar loss values, i.e. the approximate loss change $\ell(\mathbf{w}_{E,B}\!+\!\mathbf{a}^{\!-U}\!;U)\!-\!\ell(\mathbf{w}_{E,B};U)$ and the actual change $\ell(\mathbf{w}^{-U}_{E,B};U)\!-\!\ell(\mathbf{w}_{E,B};U)$ on the forgetting dataset $U$. The correlation scores range from -1 to 1, where the higher score, the more the unlearned model performs like the retrained model.
Configurations: We conduct experiments in both convex and non-convex scenarios. Specifically, we trained a multinomial Logistic Regression (LR) with total parameters $d=7850$ and a simple convolutional neural network (CNN) with total parameters $d=21840$ on MNIST dataset (Deng (2012)) for handwriting digit classification. We apply a cross-entropy loss and the inclusion of an $L^{2}$-regularization coefficient of 0.5 to ensure that the loss function of LR is strongly convex. For LR, training was performed for 15 epochs with a stepsize of 0.05 and a batch of 32. For CNN, training was carried out for 20 epochs with a stepsize of 0.05 and a batch size of 64. Given these configurations, we separately assess the distance and correlation between approximators $\mathbf{a}^{-U}_{\text{HF}}$, $\mathbf{a}^{-U}_{\text{NS}}$, $\mathbf{a}^{-U}_{\text{IJ}}$ at deletion rates in the set {1%, 5%, 10%, 15%, 20%, 25%, 30%}. Following the suggestion in Basu et al. (2021), a damping factor of 0.01 is added to the Hessian to ensure its invertibility when implementing NS and IJ. Besides, we conducted experiments with 7 random seeds to obtain average results. For the sake of clear visualizations, we use the minimum and maximum values as error bars.
The convex setting. As shown in Figure 1 (a), our approach outperforms previous works. First, by distance evaluation, the approximation error $\|\Delta_{E,B}^{-U}\|$ of the proposed method is lower than that of the previous NS and IJ works. At a deletion rate of 30%, the error is 0.2097, lower than that of NS and IJ, which have errors of 0.246554 and 0.246557, respectively. Second, in assessing the loss change for the forgotten dataset $U$, our proposed method more accurately captures the actual changes. Particularly when removing 30% of the samples, the proposed method maintains a high correlation, with Pearson and Spearman being 0.96 and 0.95, respectively. These evaluations demonstrate the efficacy of both Hessian-free and Hessian-based methods in accurately approximating the retrained model. Our proposed approach not only achieves a closer approximation to the retrained model but also exhibits lower complexity compared to previous methods. This is substantiated by the results of Application Experiments I, which provide empirical evidence supporting our theoretical findings.
The non-convex setting. As illustrated in Figure 1 (b), the proposed HF method demonstrates superior performance, exhibiting less dependency on random variations compared to NS and IJ. When 30% of the samples are removed, our method maintains a lower approximation error of 0.90, surpassing the performance of NS and IJ. Our methods achieve Spearman and Pearson correlation coefficients of 0.81 and 0.74, respectively, accurately capturing actual loss changes, in contrast to NS and IJ which are far from the actual values in non-convex settings. In Subsection D.2, we explain the volatility of the results in the prior works on non-convex setting, and provide more large-scale results, which is not achievable in NS and IJ (Sekhari et al. (2021); Suriyakumar et al. (2022); Liu et al. (2023)) due to out-of-memory to precisely computing the Hessian matrix (or its inverse).
(a) Convex Setting
(b) Non-Convex Setting
Figure 1: Verification experiments I on (a) LR and (b) CNN, respectively. The left of (a)(b) shows the distance to the retrained model, i.e. $\|\mathbf{w}_{E,B}^{-U}\!-\!\mathbf{w}_{E,B}\!\!-\!\mathbf{a}^{-U}\|$. The right of (a)(b) shows the correlation between the approximate loss change and actual loss change on the forgetting dataset.
Table 2: Application experiments I. An online scenario with 20% samples to be forgotten, where each unlearning request involves forgetting a single data point with each execution. A performance gap against Retrain is provided in ($\cdot$). We compute the total computational/storage cost for all samples.
Model Unlearning Computation Pre-Computaion Storage Test Accuracy (%)
Runtime (Sec) Speedup Runtime (Sec) (GB) Unlearned model
NS LR 5.09×$\times 10^{2}$ 1.34$\times$ 2.55×$\times 10^{3}$ 0.23 86.25 (-1.50)
CNN 5.81×$\times 10^{3}$ 0.12$\times$ 2.91×$\times 10^{4}$ 1.78 83.50 (-10.25)
IJ LR 0.23×$\times 10^{2}$ 29.64$\times$ 2.55×$\times 10^{3}$ 0.23 86.25 (-1.50)
CNN 1.91×$\times 10^{2}$ 3.63$\times$ 2.91×$\times 10^{4}$ 1.78 82.75 (-11.00)
HF LR 0.0073 93,376$\times$ 2.20×$\times 10^{2}$ 0.03 87.50 (-0.25)
CNN 0.0268 25,821$\times$ 5.34×$\times 10^{2}$ 0.08 91.50 (-2.25)
5.2 Application Experiments
We compare the cost and performance of our method with the previous certified unlearning methods. Evaluations are conducted from two perspectives: runtime and utility. Specifically, runtime focuses on the time spent pre-computing unlearning statistics and the speedup of the unlearning algorithm compared to the retraining. Moreover, we evaluate the utility by the test accuracy of unlearned model to ensure generalization guarantee. We train LR and CNN with 20% data to be forgotten, where the configuration is identical to the foregoing verification experiments I. We conduct experiments with the unlearning guarantee of $\epsilon=1,\delta=10^{-3}$ for LR and $\epsilon=100,\delta=0.1$ for CNN. We provide larger scale results in Subsection D.3, and we have not evaluated NS and IJ due to out-of-memory.
Application experiment results. Table 2 demonstrates that the proposed algorithm outperforms other baselines by a significant margin. Strikingly, the proposed algorithm outperforms benchmarks of certified unlearning regarding all metrics with millisecond-level unlearning and minimal performance degradation in both convex and non-convex settings. When models exceed the complexity of LR, precisely estimating the full Hessian becomes infeasible. This conclusion is supported by Table 2 and Mehta et al. (2022). Therefore, we demonstrate that the proposed Hessian-free perspective method exhibits great potentials for over-parameterized models compared to the Hessian-based approaches.
6 Experiments from Adversary’s Perspective
Figure 2: Membership Inference Attack I. The first plot shows the MIA-L and error bars show 95% confidence intervals; the second and third plots depict the MIA-U results; and the fourth plot illustrates the privacy-utility tradeoff. Proposed method successfully defends MIA-L and mitigates excessive privacy leakage from MIA-U without sacrificing performance when applied with appropriate noise.
To further investigate the effectiveness of data removal for auditing, we perform membership inference attack (MIA) experiments from the attacker’s perspective, including two state-of-the-art MIA: ❶ MIA against Learning (MIA-L) (LiRA MIA Carlini et al. (2022)) and ❷ MIA against Unlearning (MIA-U) (Chen et al. (2021)). Specifically, MIA-L enables an adversary to query a trained ML model to predict whether a specific example was included in the model’s training dataset; MIA-U involves querying two versions of the ML model to infer whether a target sample is part of the training set of the learned model but has been removed from the corresponding unlearned model (primarily targeting retrain algorithm). See Subsection D.4 for a detailed introduction and more experimental results.
MIA results with 5% data to be forgotten. As shown in the first plot of Figure 2, our proposed method effectively defends against MIA-L even in low-privacy scenarios, as evidenced by effective approximation of the retrained model in the verification experiment. However, MIA-U exploits discrepancies between the outputs of the model’s retrained and original versions. The verification experiment also reveals that our unlearned model’s loss change correlation coefficient is closely aligned with that of the retrained model, indicating that, like the retrained model, it is similarly vulnerable to such attacks. As depicted in the second and third plots, MIA-U shows high AUC values for the retrained method (light red line) and our method (red line) without perturbation ($\epsilon=\infty$). Nevertheless, adding appropriate noise to the model can mitigate MIA-U. As demonstrated in the fourth plot, introducing an appropriate noise level ($\epsilon>2$ for LR and $\epsilon>120$ for CNN) effectively reduces the excessive privacy leakage without compromising performance. Besides, further increasing the model perturbation to tradeoff privacy-utility does not significantly reduce the effectiveness of such attacks. To better defend MIA-U, we will consider measures like noised SGD in future work.
7 Conclusion
We proposed a novel Hessian-free certified unlearning method. Our analysis relies on an affine stochastic recursion to characterize the model update discrepancy between the learned and retrained models, which does not depend on inverting Hessian and model optimality. We provide results of unlearning, generalization, deletion capacity guarantees, and time/memory complexity that are superior to state-of-the-art methods. We then develop an algorithm that achieves near-instantaneous unlearning. We show that the proposed method can reduce the precomputation time, storage, and unlearning time to $\mathcal{O}(n^{2}d)$, $\mathcal{O}(nd)$, $\mathcal{O}(d)$, respectively. Experimental results validate that our work benefits certified unlearning with millisecond-level unlearning time and improved test accuracy.
Acknowledgments
This work was supported in part by NSF grants CMMI-2024774 and ECCS-2030251, and in part by the National Natural Science Foundation of China under Grant 62202427.
Appendix A An Expanded Version of The Related Work
Machine Unlearning, the terminology can be traced back to the paper of Cao & Yang (2015), and their work defined the notion of "the completeness of forgetting". There are two main categories based on the completeness of forgetting: Exact Unlearning and Approximate Unlearning.
Exact Unlearning aims to accelerate the speed of retraining and completely eliminate the influence of forgotten data on learned model. The majority of existing Exact Unlearning methods are derived from the work of Bourtoule et al. (2021), which proposed the notable SISA framework, grounded in ensemble learning techniques. Specifically, the training dataset is partitioned into mutually disjoint shards, and then sub-models are trained on these shards. Upon unlearning, SISA only needs to retrain the sub-model correlated with the relevant data shard and then make the final prediction by assembling the knowledge of each sub-model, which significantly lowers retraining computational costs. Under the SISA framework, other Exact Unlearning methods have also been proposed, e.g., Chen et al. (2022b) extended SISA to graph learning by partitioning the original training graph into disjoint shards, parallelly training a set of shard models, and learning an optimal importance score for each shard model. RecEraser (Chen et al. (2022a)) extended SISA to recommendation systems. KNOT (Su & Li (2023)) adopted the SISA for client-level asynchronous federated unlearning during training and introduced a clustered mechanism that divides clients into multiple clusters. Other methods, such as DaRE (Brophy & Lowd (2021)) and HedgeCut (Schelter et al. (2021)), addressed the issue that SISA cannot be applied to decision trees and proposed exact unlearning solutions; and Ginart et al. (2019) aimed to achieve forgetting in K-means clustering. Furthermore, ARCANE (Yan et al. (2022)) introduced a class-based data partitioning strategy to enhance the performance of SISA.
However, the SISA framework adopts a structure that completely separates the data, resembling ensemble learning approaches where predictions or models are aggregated only at the end. This separation restricts the model training process from fully leveraging the entirety of the segmented datasets and also hampers its extensibility, leading to lower accuracy and performance compared to a model trained on the entire unsegmented dataset. In an experimental evaluation conducted by Koch & Soll (2023), it was found that simply down-sampling the data set to achieve a more balanced single shard outperforms SISA in terms of unlearning speed and performance.
Approximate Unlearning thus seeks to minimize the impact of forgetting data to an acceptable level to trade off computational efficiency, reduced storage costs, and flexibility. There are many approaches to approximate unlearning, such as Kurmanji et al. (2024) developed a distillation-based method; Halimi et al. (2022) adopted a gradient ascent approach on forgotten data to achieve forgetting. Among these, second-order methods have been widely studied, with most of these methods (e.g., Sekhari et al. (2021); Suriyakumar et al. (2022); Liu et al. (2023); Chien et al. (2022); Wu et al. (2023); Mehta et al. (2022); Warnecke et al. (2023); Wu et al. (2022)) inspired by influence function studies (Koh & Liang (2017); Basu et al. (2021; 2020); Hara et al. (2019)), aiming to extract information from the Hessian of loss function and achieve forgetting through limited model updates. In particular, inspired by differential privacy (Dwork (2006)), Guo et al. (2020) introduced certified unlearning $(\epsilon,\delta)$, anticipating ensuring that the output distribution of an unlearning algorithm is indistinguishable from that of retraining. Moreover, due to costly Hessian operations, Fisher-based methods (Golatkar et al. (2020a; b; 2021)) have emerged to approximate Hessian. Our paper concentrates on the Hessian-based approach, yet obviates the need for Hessian inversion and explicit calculation of the Hessian.
Among these Hessian-based certified unlearning methods, Sekhari et al. (2021); Suriyakumar et al. (2022); Liu et al. (2023) established the theoretical foundation for Hessian-based methods and provided the generalization result in terms of the population risk, as well as derived the deletion capacity guarantee. In the definition of certified unlearning, their works align with our method, which lies in pre-computing statistics of data that extract second-order information from the Hessian and using these ‘recollections’ to forget data. However, existing second-order methods still require expensive matrix computations and can only perform forgetting for optimal models under convexity assumption. Although Zhang et al. (2024) offered certified unlearning for deep model through adding sufficient large regularization to guarantee local strong convexity, but the optimality problem remains unsolved, particularly due to the non-convexity of modern deep models, making it impractical to find the empirical risk minimizer. Even for convex models, in most non-concurrent training scenarios, such as online learning or continual learning, the model cannot converge to the optimal solution.
Appendix B Discussions
B.1 Limitations and Future Works
Limitations. Although we have made the above contributions, there are some significant limitations:
① Our proposed algorithm necessitates $\mathcal{O}(n^{2}d)$ precomputation and $\mathcal{O}(nd)$ storage, which remains costly, particularly in scenarios involving extensive data and larger-scale models. Consequently, conducting experiments concurrently with both large-scale datasets and models is anticipated to significantly extend the time required for precomputing unlearning statistics.
② As demonstrated by the results in Theorem 4, it is important to recognize that the proposed method is highly sensitive to the the stability of the training process or well-conditioned Hessian on non-convex setting. The addition of a large explicit $L^{2}$-regularizer to the objective or utilizing small step sizes can ensure a meaningful bound. However, we are not aware of any popular non-convex machine learning problem that can satisfy all the conditions of Theorem 4. If the learning process is unstable, unlearning will introduce significant error. For instance, in Subsection E.1, employing an aggressive step size ($\eta>0.1$) can lead to a substantial increase in approximation error.
③ While we propose a new method that does not rely on convexity, our analysis builds upon the theoretical framework of previous works for comparison purposes and much of our theoretical results rely on convexity—except for our unlearning-related conclusions. Providing a nonconvex theoretical framework is meaningful, and we will explore potential extensions in further research.
Future work. Therefore, to solve the problems mentioned above, we briefly propose some solutions.
① Specifically, our proposed framework offers the feasibility to unlearn an arbitrary subset of the entire datasets, allowing a trade-off that further reduces the precomputation and storage complexity, for instance, consider a typical scenario at the user level, where there are $k$ users ($k<<n$). By focusing on the unlearning of a single user’s dataset or a subset of users’ datasets, we can maintain $k$ vectors for $k$ users’ dataset instead of maintaining $n$ vectors for $n$ all data points. Therefore, we can further reduce precomputation and storage to $\mathcal{O}(knd)$ and $\mathcal{O}(kd)$, respectively.
② Moreover, based on our analysis in Theorem 4 and Subsection E.1, the main source of the approximation error is $\rho$, which mainly consists of two components: the step size $\eta$ and the maximum eigenvalue $\lambda_{1}$ of Hessian. Reducing the step size is straightforward method, but we cannot expect every training process to have a sufficiently small step size. Therefore, reducing $\lambda_{1}$ during optimization, will not only improve the generalization ability of the learned model but also reduce the approximation error in the unlearning stage. Possible approaches are through stepsize warm-up strategies (Gilmer et al. (2022)) or sharpness-aware minimization (Wen et al. (2023)) to effectively reduce $\lambda_{1}$.
B.2 Deletion Capacity Guarantee
We analyze the deletion capacity of our method based on the analysis of Sekhari et al. (2021), which formalizes how many samples can be deleted from the model parameterized by the original empirical risk minimizer while maintaining reasonable guarantees on the test loss. We establish the relationship between generalization performance and the amount of deleted samples through deletion capacity, and our analysis indicates that our proposed method outperforms previous works in both generalization and deletion capacity. The lower bound on deletion capacity can be directly derived from the generalization guarantee (Theorem 6). Below, we first formalize the deletion capacity.
Definition 2 (Definition of Deletion Capacity (Sekhari et al. (2021))).
Let $\epsilon,\delta\geq 0$. Let $S$ be a dataset of size $n$ drawn i.i.d. from $\mathcal{D}$. For a pair of learning and unlearning algorithms ${\Omega},\bar{\Omega}$ that are $(\epsilon,\delta)$-certified unlearning, the deletion capacity $m_{\epsilon,\delta}^{\Omega,\bar{\Omega}}(d,n)$ is defined as the maximum number of samples $U$ to be forgotten, while still ensuring an excess population risk is $\gamma$. Specifically,